# Fractionally modulated discrete Carleson’s Theorem and pointwise Ergodic Theorems along certain curves

### v1

@@ 5. Reducing the proof of Theorem 1 to the continuous case
$\Big{|}\sum_{m\in I^{1}_{x,y}}e\left(\lambda_{x}|x-m|^{c}-\lambda_{y}|y-m|^{c}\right)\psi^{-}_{j}(x-m)\psi_{j}^{-}(y-m)\Big{|}\lesssim 2^{\frac{\nu^{\prime}}{2}j+\frac{r}{2}-2j}+2^{-\frac{r}{2}-j+\frac{\nu^{\prime}}{2}j}$
$=2^{-j}\big{(}2^{\frac{\nu^{\prime}}{2}j+\frac{r}{2}-j}+2^{-\frac{r}{2}+\frac{\nu^{\prime}}{2}j}\big{)}\text{,}$
and note that $\frac{\nu^{\prime}}{2}j+\frac{r}{2}-j<\frac{\nu^{\prime}}{2}j-j+\frac{c+2\delta_{1}}{2}j+1=\frac{\nu^{\prime}+c-2+2\delta_{1}}{2}j+1$ and since $\nu^{\prime}\leq(2-c)/2$ and $\delta_{1}<(2-c)/4$ we get $\frac{\nu^{\prime}+c-2+2\delta_{1}}{2}<0$. We also have that $-\frac{r}{2}+\frac{\nu^{\prime}}{2}j\leq\frac{\nu^{\prime}}{2}j-\frac{\nu}{2}j+1=\frac{\nu^{\prime}-\nu}{2}j+1$, and since $\nu^{\prime}<\nu/2$ we get that $\frac{\nu^{\prime}-\nu}{2}<0$. The estimate is complete and the same reasoning applies for $I^{2}_{x,y}$. ∎
@@
(5.15) $|T_{\tau,j}T_{\tau,j}^{*}f(x)|\lesssim(2^{-\rho_{4}j}+2^{-\delta_{2}j})Mf(x)\text{,}$
and thus we get $\|T_{\tau,j}\|_{\ell^{2}\to\ell^{2}}\lesssim 2^{-\rho_{3}j}$ for $\rho_{3}=\min\big{\{}\frac{\delta_{2}}{2},\frac{\rho_{4}}{2}\big{\}}>0$, and the proof is complete. ∎
@@ 6. Concluding the Proof of Theorem 1
This section is devoted to establishing the estimate (6.1) which we do in the following proposition, concluding the proof of Theorem 1. Similarly to previously, we have fixed $c\in(1,2)$, $\varepsilon\in\big{(}0,\min\big{\{}\frac{1}{4},2-c\big{\}}\big{)}$ and $\nu\in(0,\varepsilon/10)$ and all implied constants may depend on them.
@@
Let $\mathsf{i}\in\{0,1\}$, $p\in(1,\infty)$. There exists a positive constant $C=C(p,c,\varepsilon,\nu)$ such that for all $f\in L^{p}(\mathbb{R})$ we have
@@
$\mathcal{C}^{\mathsf{i}}_{\varepsilon}f(x)=\sup_{\lambda\in\mathbb{R}}\bigg{|}\int_{\mathbb{R}}f(x-t)\frac{e(\lambda[t]_{\mathsf{i}}^{\varepsilon})}{t}{\rm d}t\bigg{|}=\sup_{\lambda\in\mathbb{R}}\Big{|}\mathcal{F}_{\mathbb{R}}^{-1}\Big{[}\mathcal{F}_{\mathbb{R}}[f](\xi)\;\mathsf{p.v.}\int_{\mathbb{R}}\frac{e(\lambda[t]_{\mathsf{i}}^{\varepsilon}-\xi t)}{t}{\rm d}t\Big{]}(x)\Big{|}\text{.}$
@@
where $j(\lambda,\ell)$ is the unique integer such that $21\leq\lambda^{\frac{1}{c}}2^{j(\lambda,\ell)-\ell}<$ 2. Therefore we may write
@@
(6.4) $S_{1}\leq\big{\|}\mathcal{C}_{c}^{\mathsf{i}}\big{[}f*\mathcal{F}_{\mathbb{R}}^{-1}[\eta_{1}]\big{]}\big{\|}_{L^{p}(\mathbb{R})}\lesssim\|f*\mathcal{F}_{\mathbb{R}}^{-1}[\eta_{1}]\|_{L^{p}(\mathbb{R})}\lesssim_{\eta_{1}}\|f\|_{L^{p}(\mathbb{R})}\text{,}$
where we have used Theorem 5. The estimate for $S_{1}$ is complete.
@@
Let $\mathsf{i}\in\{0,1\}$. For every $p\in(1,\infty)$ there exist positive constants $C=C(c,p)$ and $\beta=\beta(c,p)$ such that for every $\ell\in\mathbb{Z}$ we have
@@
(6.5) $S_{2}\leq\sum_{\ell\geq\frac{\nu}{c}-1}\bigg{\|}\sup_{\lambda\in[-1/2,1/2)}\Big{|}\mathcal{F}^{-1}_{\mathbb{R}}\Big{[}\mathcal{F}_{\mathbb{R}}[f](\cdot)\bm{1}_{[1,2)}(\lambda^{\frac{1}{c}}2^{j-\ell})H^{\mathsf{i}}_{j(\lambda,\ell)}(\cdot,\lambda))\bm{1}_{X_{j(\lambda,\ell)}^{c}}(\lambda)\Big{]}\Big{|}\bigg{\|}_{L^{p}(\mathbb{R})}\\ \leq\sum_{\ell\geq\frac{\nu}{c}-1}\Big{\|}\sup_{j\in\mathbb{Z}}\sup_{|\lambda|^{\frac{1}{c}}2^{j}\simeq 2^{\ell}}\Big{|}\mathcal{F}_{\mathbb{R}}^{-1}\big{[}\mathcal{F}_{\mathbb{R}}(f)(\cdot)H_{j}^{\mathsf{i}}(\cdot,\lambda)\big{]}\Big{|}\Big{\|}_{L^{p}(\mathbb{R})}\\ \lesssim\sum_{\ell\geq\frac{\nu}{c}-1}2^{-\beta\ell}\|f\|_{L^{p}(\mathbb{R})}\lesssim_{p}\|f\|_{L^{p}(\mathbb{R})}\text{,}$
and the estimate for $S_{2}$ is complete. We now provide a proof of the above lemma.
@@
With an argument analogous to the one given in Section 5, one sees that it suffices to establish that the $L^{2}$ norm of the following operator
@@
is controlled by $2^{-\beta\ell}$ for some $\beta>0$, and for any choice of measurable functions $\lambda_{x}\colon\mathbb{Z}\to[0,\infty)$ and $j_{x}\colon\mathbb{Z}\to\mathbb{Z}$ with the property that $\lambda_{x}^{\frac{1}{c}}2^{j_{x}}\simeq 2^{\ell}$, where the implied constants do not depend on the choice functions. The corresponding result on $L^{p}$ follows from interpolation with identical considerations with the ones of Section 5. To that end we prove a pointwise bound on the kernel of the associated $TT^{*}$ operator defined by
@@
More precisely, we prove that for a small parameter $\delta\in(0,\frac{c}{100})$ we have the following pointwise bound
(6.7) $\left|K_{\mathsf{i},\tau,\ell}(x,y)\right|\lesssim 2^{-\max\left\{j_{x},j_{y}\right\}}\bm{1}_{|x-y|\lesssim 2^{\max\left\{j_{x},j_{y}\right\}-\ell\delta}}+2^{-\max\left\{j_{x},j_{y}\right\}-\ell\rho}\bm{1}_{2^{\max\left\{j_{x},j_{y}\right\}-\ell\delta}\lesssim|x-y|\lesssim 2^{\max\left\{j_{x},j_{y}\right\}}}\text{,}$
for some $\rho>0$. This establishes the desired $L^{2}$-bound with an argument as in (5.15). We have that
@@
and thus if $|x-y|\lesssim 2^{\max\left\{j_{x},j_{y}\right\}-\ell\delta}$ we have that (6.7) is satisfied. Therefore, we are going to assume that $|x-y|\gtrsim 2^{\max\left\{j_{x},j_{y}\right\}-\ell\delta}$. Without loss of generality let $x\leq y.$ We are also going to denote by $g$ the phase of the oscillatory integral we are interested in. Let us mention that all $t\in[x,y]$ may be discarded in the integration in (6.6) since $x-t$, $y-t$ will have different sign and the integrand will be zero. We begin with the case $|j_{x}-j_{y}|\geq 100$. We have
$\begin{split}&\partial_{t}g(t)=\varepsilon_{t,x}\lambda_{x}|x-t|^{c-1}-\varepsilon_{t,y}\lambda_{y}|y-t|^{c-1},\quad\text{with }|\varepsilon_{t,x}|=|\varepsilon_{t,y}|=1\\ &\Rightarrow|\partial_{t}g|\gtrsim 2^{c\ell}2^{-\min\left\{j_{x},j_{y}\right\}}\end{split}$
@@
If $\varepsilon(x,y)\leq 2^{-\delta\ell-100}$, then for $t\in I_{x,y}\setminus[x,y]$ we have
$|\partial_{t}g(t)|=c\left|\lambda_{x}|t-x|^{c-1}-\lambda_{y}|t-y|^{c-1}\right|\simeq\xi^{c-2}\left|\lambda_{x}^{\frac{1}{c-1}}|t-x|-\lambda_{y}^{\frac{1}{c-1}}|t-y|\right|\text{,}$
for some $\xi$ between $\lambda_{y}^{\frac{1}{c-1}}|t-y|$ and $\lambda_{x}^{\frac{1}{c-1}}|t-x|$. We also used the fact that in $I_{x,y}\setminus[x,y]$, $t-x$ and $t-y$ have the same sign. We also have
$\begin{split}|\partial_{t}g(t)|&\gtrsim\xi^{c-2}|\lambda_{x}^{\frac{1}{c-1}}(t-x)-\lambda_{y}^{\frac{1}{c-1}}(t-y)|\gtrsim\lambda_{x}2^{cj_{x}}2^{-2j_{x}}|t-x-\left(\frac{\lambda_{y}}{\lambda_{x}}\right)^{\frac{1}{c-1}}(t-y)|\\ &\gtrsim 2^{c\ell-2j_{x}}|t-x-(1+\varepsilon(x,y))(t-y)|\gtrsim 2^{c\ell-2j_{x}}2^{j_{x}-\delta\ell}\gtrsim 2^{(c-\delta)\ell-j_{x}}\text{.}\end{split}$
@@
$\left|\int_{I_{x,y}\setminus[x,y]}e(\lambda_{x}[x-t]_{\mathsf{i}}^{c}-\lambda_{y}[y-t]_{\mathsf{i}}^{c})\psi^{\tau}_{j_{x}}(x-t)\psi^{\tau}_{j_{y}}(y-t){\rm d}t\right|\lesssim\frac{2^{-2j_{x}}}{2^{-j_{x}+(c-\delta)\ell}}=2^{-(c-\delta)\ell-j_{x}}\text{,}$
@@
Finally, if $\varepsilon(x,y)>2^{-\delta\ell-100}$ by recasting the above calculation we have that
(6.8) $E_{x,y}\coloneqq\left\{t\in I_{x,y}\setminus[x,y]:|\partial_{t}g(t)|\leq 2^{-j_{x}+\delta\ell}\right\}\\ \subseteq\left\{t\in I_{x,y}\setminus[x,y]:2^{c\ell-2j_{x}}|y-x+\varepsilon(x,y)(t-y)|\lesssim 2^{-j_{x}+\delta\ell}\right\}\\ \subseteq\left\{t\in I_{x,y}\setminus[x,y]:|y-x+\varepsilon(x,y)(t-y)|\lesssim 2^{j_{x}+(-c+\delta)\ell}\right\}\text{,}$
which yields that $|E_{x,y}|\lesssim 2^{j_{x}+(-c+2\delta)\ell}$. Therefore we can split the oscillatory integral on $I_{x,y}\setminus[x,y]$ as follows
$\begin{split}\left|\int_{I_{x,y}\setminus[x,y]}e(\lambda_{x}[x-t]_{\mathsf{i}}^{c}-\lambda_{y}[y-t]_{\mathsf{i}}^{c})\psi^{\tau}_{j_{x}}(x-t)\psi^{\tau}_{j_{y}}(y-t){\rm d}t\right|&\leq\left|\int_{E_{x,y}}\cdots\right|+\left|\int_{I_{x,y}\setminus E_{x,y}}\cdots\right|\\ &\lesssim 2^{-2j_{x}+j_{x}+(-c+\delta)\ell}+\frac{2^{-2j_{x}}}{2^{-j_{x}+\delta\ell}}\\ &\lesssim 2^{-j_{x}+(-c+2\delta)\ell}+2^{-j_{x}-\delta\ell}\text{.}\end{split}$
Since $-c+2\delta<0$ and $\delta>0$ the proof is complete. ∎

### v2

@@ 5. Reducing the proof of Theorem 1 to the continuous case
(5.30) $\Big|\sum_{m\in I^{1}_{x,y}}e\left(\lambda_{x}|x-m|^{c}-\lambda_{y}|y-m|^{c}\right)\psi^{-}_{j}(x-m)\psi_{j}^{-}(y-m)\Big|\\ \lesssim 2^{\frac{\nu^{\prime}}{2}j+\frac{r}{2}-2j}+\big((2-c)(c-1)\big)^{-\frac{1}{2}}2^{-\frac{r}{2}-j+\frac{\nu^{\prime}}{2}j}\lesssim\big((2-c)(c-1)\big)^{-\frac{1}{2}}2^{-j}\Big(2^{\frac{\nu^{\prime}}{2}j+\frac{r}{2}-j}+2^{-\frac{r}{2}+\frac{\nu^{\prime}}{2}j}\Big)\\ \lesssim\big((2-c)(c-1)\big)^{-\frac{1}{2}}2^{-j(1+\frac{\varepsilon}{64})}\text{,}$
where for the last estimate we took into account that $r\in\big[\nu j-2,(c+2\varepsilon)j+2\big]$ as well as the fact that $\nu^{\prime}=\frac{3\varepsilon}{32}$, $\nu=\frac{\varepsilon}{8}$, $\varepsilon\coloneqq\frac{2-c}{4}$ and thus
$\frac{\nu^{\prime}}{2}j+\frac{r}{2}-j<\frac{\nu^{\prime}}{2}j-j+\frac{c+2\varepsilon}{2}j+1=\frac{\nu^{\prime}+c-2+2\varepsilon}{2}j+1=\frac{\frac{3\varepsilon}{32}-4\varepsilon+2\varepsilon}{2}j+1\leq-\frac{\varepsilon}{64}j+1$
and
$\frac{r}{2}+\frac{\nu^{\prime}}{2}j\leq\frac{\nu^{\prime}}{2}j-\frac{\nu}{2}j+1=\frac{\frac{3\varepsilon}{32}-\frac{4\varepsilon}{32}}{2}j+1\leq-\frac{\varepsilon}{64}j+1\text{.}$
The estimate is complete and the same reasoning applies for $I^{2}_{x,y}$. ∎
@@
(5.31) $|T_{\tau,j}T_{\tau,j}^{*}f(x)|\lesssim\big((2-c)(c-1)\big)^{-\frac{1}{2}}(2^{-\rho j}+2^{-\sigma j})Mf(x)\lesssim\big((2-c)(c-1)\big)^{-\frac{1}{2}}2^{-\rho j}Mf(x)\text{,}$
and thus we get $\|T_{\tau,j}\|_{\ell^{2}\to\ell^{2}}\lesssim\big((2-c)(c-1)\big)^{-\frac{1}{4}}2^{-\frac{\rho}{2}j}$, establishing (5.20) and concluding the proof. ∎
@@ 6. Concluding the Proof of Theorem 1
This section is devoted to establishing the estimate (6.1) which we do in the following proposition, concluding the proof of Theorem 1.
@@
Let $\mathsf{i}\in\{0,1\}$, $p\in(1,\infty)$. There exists a positive constant $C=C(p,c)$ such that for all $f\in L^{p}(\mathbb{R})$ we have
@@
(6.2) $\mathcal{C}^{\mathsf{i}}_{\varepsilon}f(x)=\sup_{\lambda\in\mathbb{R}}\bigg|\int_{\mathbb{R}}f(x-t)\frac{e(\lambda[t]_{\mathsf{i}}^{\varepsilon})}{t}{\rm d}t\bigg|=\sup_{\lambda\in\mathbb{R}}\Big|\mathcal{F}_{\mathbb{R}}^{-1}\Big[\mathcal{F}_{\mathbb{R}}[f](\xi)\;\mathsf{p.v.}\int_{\mathbb{R}}\frac{e(\lambda[t]_{\mathsf{i}}^{\varepsilon}-\xi t)}{t}{\rm d}t\Big](x)\Big|\text{.}$
@@
We briefly mention that for $\varepsilon\in(1,2)$ we have $C_{2,\varepsilon,\mathsf{i}}\lesssim(c-1)^{-\frac{1}{2}}$, where the implied constant is absolute. This is shown in Lemma 6.3.
@@
where $j(\lambda,\ell)$ is the unique integer such that $1\leq\lambda^{\frac{1}{c}}2^{j(\lambda,\ell)-\ell}<2$. Therefore we may write
@@
(6.5) $S_{1}\leq\big\|\mathcal{C}_{c}^{\mathsf{i}}\big[f*\mathcal{F}_{\mathbb{R}}^{-1}[\eta_{1}]\big]\big\|_{L^{p}(\mathbb{R})}\lesssim_{c,p}\|f*\mathcal{F}_{\mathbb{R}}^{-1}[\eta_{1}]\|_{L^{p}(\mathbb{R})}\lesssim_{\eta_{1}}\|f\|_{L^{p}(\mathbb{R})}\text{,}$
where we have used Theorem 5. The estimate for $S_{1}$ is complete, and we note that for the case $p=2$, we obtain $S_{1}\lesssim(c-1)^{-\frac{1}{2}}\|f\|_{L^{2}(\mathbb{R})}$, see our comment after Theorem 5 and Lemma 6.3.
@@
Let $\mathsf{i}\in\{0,1\}$. For every $p\in(1,\infty)$ there exist positive constants $C=C(c,p)$ and $\beta=\beta(c,p)$ such that for every $\ell\in\mathbb{N}$ we have
@@
Moreover, for $p=2$, one may choose $C\lesssim(c-1)^{-\frac{1}{2}}$ and $\beta=\frac{1}{6}$.
@@
(6.6) $S_{2}\leq\sum_{\ell\geq\frac{\nu}{c}-1}\bigg\|\sup_{\lambda\in[-1/2,1/2)}\Big|\mathcal{F}^{-1}_{\mathbb{R}}\Big[\mathcal{F}_{\mathbb{R}}[f](\cdot)\bm{1}_{[1,2)}(\lambda^{\frac{1}{c}}2^{j-\ell})H^{\mathsf{i}}_{j(\lambda,\ell)}(\cdot,\lambda))\bm{1}_{X_{j(\lambda,\ell)}^{c}}(\lambda)\Big]\Big|\bigg\|_{L^{p}(\mathbb{R})}\\ \leq\sum_{\ell\geq\frac{\nu}{c}-1}\Big\|\sup_{j\in\mathbb{Z}}\sup_{|\lambda|^{\frac{1}{c}}2^{j}\simeq 2^{\ell}}\Big|\mathcal{F}_{\mathbb{R}}^{-1}\big[\mathcal{F}_{\mathbb{R}}(f)(\cdot)H_{j}^{\mathsf{i}}(\cdot,\lambda)\big]\Big|\Big\|_{L^{p}(\mathbb{R})}\\ \lesssim_{c,p}\sum_{\ell\geq\frac{\nu}{c}-1}2^{-\beta\ell}\|f\|_{L^{p}(\mathbb{R})}\lesssim_{p}\|f\|_{L^{p}(\mathbb{R})}\text{,}$
and the estimate for $S_{2}$ is complete. For the case $p=2$ we additionally note that by taking into account the second assertion of Lemma 6.2 we get
(6.7) $S_{2}\lesssim(c-1)^{-\frac{1}{2}}\Big(\sum_{\ell\geq\frac{\nu}{c}-1}2^{-\frac{\ell}{6}}\Big)\|f\|_{L^{2}(\mathbb{R})}\leq(c-1)^{-\frac{1}{2}}\Big(\sum_{\ell\geq-1}2^{-\frac{\ell}{6}}\Big)\|f\|_{L^{2}(\mathbb{R})}\lesssim(c-1)^{-\frac{1}{2}}\|f\|_{L^{2}(\mathbb{R})}\text{.}$
We now provide a proof of the above lemma.
@@
With an argument analogous to the one given in Section 5, one sees that it suffices to establish the second assertion of the lemma and specifically we prove that the $L^{2}$ norm of the following operator
@@
is controlled by $\lesssim(c-1)^{-\frac{1}{2}}2^{-\frac{\ell}{6}}$ for any choice of measurable functions $\lambda_{x}\colon\mathbb{Z}\to[0,\infty)$ and $j_{x}\colon\mathbb{Z}\to\mathbb{Z}$ with the property that $\lambda_{x}^{\frac{1}{c}}2^{j_{x}}\simeq 2^{\ell}$, where the implied constants are absolute. The corresponding result on $L^{p}$ follows from interpolation with identical considerations with the ones of Section 5. To that end we prove a pointwise bound on the kernel of the associated $TT^{*}$ operator defined by
@@
More precisely, we prove that for $\sigma\coloneqq\frac{1}{3}$ we have the following pointwise bound
(6.9) $\left|K_{\mathsf{i},\tau,\ell}(x,y)\right|\\ \lesssim(c-1)^{-1}2^{-\max\left\{j_{x},j_{y}\right\}-\sigma\ell}\bm{1}_{2^{\max\left\{j_{x},j_{y}\right\}-\sigma\ell}\lesssim|x-y|\lesssim 2^{\max\left\{j_{x},j_{y}\right\}}}+2^{-\max\left\{j_{x},j_{y}\right\}}\bm{1}_{|x-y|\lesssim 2^{\max\left\{j_{x},j_{y}\right\}-\sigma\ell}}\text{.}$
This establishes the desired $L^{2}$-bound with an argument as in (5.31). We have that
@@
and thus if $|x-y|\lesssim 2^{\max\left\{j_{x},j_{y}\right\}-\sigma\ell}$ we have that (6.9) is satisfied. Therefore, we are going to assume that $|x-y|\gtrsim 2^{\max\left\{j_{x},j_{y}\right\}-\sigma\ell}$. Without loss of generality let $x\leq y.$ We are also going to denote by $g$ the phase of the oscillatory integral we are interested in. Let us mention that all $t\in[x,y]$ may be discarded in the integration in (6.8) since xtx-t, yty-t will have different signs and the integrand will be zero. We begin with the case $|j_{x}-j_{y}|\geq 100$. We have
$\begin{split}&\partial_{t}g(t)=\varepsilon_{t,x}\lambda_{x}c|x-t|^{c-1}-\varepsilon_{t,y}\lambda_{y}c|y-t|^{c-1},\quad\text{with }|\varepsilon_{t,x}|=|\varepsilon_{t,y}|=1\\ &\Rightarrow|\partial_{t}g|\gtrsim 2^{c\ell}2^{-\min\left\{j_{x},j_{y}\right\}}\end{split}$
@@
If $\varepsilon(x,y)\leq 2^{-\frac{\ell}{3}-100}$, then for $t\in I_{x,y}\setminus[x,y]$ we have
$|\partial_{t}g(t)|=c\left|\lambda_{x}|t-x|^{c-1}-\lambda_{y}|t-y|^{c-1}\right|\simeq c(c-1)\xi^{c-2}\left|\lambda_{x}^{\frac{1}{c-1}}|t-x|-\lambda_{y}^{\frac{1}{c-1}}|t-y|\right|\text{,}$
for some $\xi$ between $\lambda_{y}^{\frac{1}{c-1}}|t-y|$ and $\lambda_{x}^{\frac{1}{c-1}}|t-x|$. We also used the fact that in $I_{x,y}\setminus[x,y]$, txt-x and tyt-y have the same sign. We also have
$\begin{split}|\partial_{t}g(t)|&\gtrsim(c-1)\xi^{c-2}|\lambda_{x}^{\frac{1}{c-1}}(t-x)-\lambda_{y}^{\frac{1}{c-1}}(t-y)|\gtrsim(c-1)\lambda_{x}2^{cj_{x}}2^{-2j_{x}}|t-x-\left(\frac{\lambda_{y}}{\lambda_{x}}\right)^{\frac{1}{c-1}}(t-y)|\\ &\gtrsim(c-1)2^{c\ell-2j_{x}}|t-x-(1+\varepsilon(x,y))(t-y)|\gtrsim(c-1)2^{c\ell-2j_{x}}2^{j_{x}-\frac{\ell}{3}}\gtrsim(c-1)2^{(c-\frac{1}{3})\ell-j_{x}}\text{.}\end{split}$
@@
(6.10) $\left|\int_{I_{x,y}\setminus[x,y]}e(\lambda_{x}[x-t]_{\mathsf{i}}^{c}-\lambda_{y}[y-t]_{\mathsf{i}}^{c})\psi^{\tau}_{j_{x}}(x-t)\psi^{\tau}_{j_{y}}(y-t){\rm d}t\right|\\ \lesssim(c-1)^{-1}\frac{2^{-2j_{x}}}{2^{-j_{x}+(c-\frac{1}{3})\ell}}=(c-1)^{-1}2^{-(c-\frac{1}{3})\ell-j_{x}}\leq(c-1)^{-1}2^{-\frac{1}{3}\ell-j_{x}}\text{,}$
@@
Finally, if $\varepsilon(x,y)>2^{-\frac{\ell}{3}-100}$ by recasting the above calculation we have that
(6.11) $E_{x,y}\coloneqq\left\{t\in I_{x,y}\setminus[x,y]:|\partial_{t}g(t)|\leq 2^{-j_{x}+\frac{\ell}{3}}\right\}\\ \subseteq\left\{t\in I_{x,y}\setminus[x,y]:2^{c\ell-2j_{x}}|y-x+\varepsilon(x,y)(t-y)|\lesssim\big(c(c-1)\big)^{-1}2^{-j_{x}+\frac{\ell}{3}}\right\}\\ \subseteq\left\{t\in I_{x,y}\setminus[x,y]:|y-x+\varepsilon(x,y)(t-y)|\lesssim(c-1)^{-1}2^{j_{x}+(-c+\frac{1}{3})\ell}\right\}\text{,}$
which yields that $|E_{x,y}|\lesssim(c-1)^{-1}2^{j_{x}+(-c+\frac{2}{3})\ell}\lesssim(c-1)^{-1}2^{j_{x}-\frac{1}{3}\ell}$. Therefore we can split the oscillatory integral on $I_{x,y}\setminus[x,y]$ as follows
(6.12) $\left|\int_{I_{x,y}\setminus[x,y]}e(\lambda_{x}[x-t]_{\mathsf{i}}^{c}-\lambda_{y}[y-t]_{\mathsf{i}}^{c})\psi^{\tau}_{j_{x}}(x-t)\psi^{\tau}_{j_{y}}(y-t){\rm d}t\right|\leq\left|\int_{E_{x,y}}\cdots\right|+\left|\int_{I_{x,y}\setminus E_{x,y}}\cdots\right|\\ \lesssim(c-1)^{-1}2^{-2j_{x}+j_{x}-\frac{\ell}{3}}+(c-1)^{-1}\frac{2^{-2j_{x}}}{2^{-j_{x}+\frac{\ell}{3}}}\lesssim(c-1)^{-1}2^{-j_{x}-\frac{1}{3}\ell}\text{,}\quad\text{as desired.}$
The proof is complete. ∎
Let us note that with this lemma essentially yields Theorem 5 on $L^{2}$ for $c\in(1,2)$. Since we would like to keep track of the dependencies of constants with respect to $c\in(1,2)$ on $L^{2}$ we quickly showcase how this is done while establishing the bound $C_{2,\varepsilon,\mathsf{i}}\lesssim(c-1)^{-\frac{1}{2}}$ we claimed earlier, see Theorem 5 above.
Lemma 6.3.
For every $c\in(1,2)$ and $\mathsf{i}\in\{0,1\}$ we have
$\big\|\mathcal{C}^{\mathsf{i}}_{c}f\big\|_{L^{2}(\mathbb{R})}\lesssim(c-1)^{-\frac{1}{2}}\|f\|_{L^{2}(\mathbb{R})}\text{,}$
where the operator $\mathcal{C}^{\mathsf{i}}_{c}$ is defined in (6.2) and the implicit constant is absolute.
Proof.
It suffices to establish the estimate for smooth compactly supported functions and we note that
(6.13) $\int_{\mathbb{R}}f(x-t)\frac{e(\lambda[t]_{\mathsf{i}}^{c})}{t}{\rm d}t=\sum_{j\in{\mathbb{Z}}}\int_{\mathbb{R}}f(x-t)e(\lambda[t]_{\mathsf{i}}^{c})\psi_{j}(t){\rm d}t\\ =\sum_{j\in{\mathbb{Z}}}\sum_{r\in{\mathbb{Z}}}\bm{1}_{[2^{r},2^{r+1})}(|\lambda|^{\frac{1}{c}}2^{j})\int_{\mathbb{R}}f(x-t)e(\lambda[t]_{\mathsf{i}}^{c})\psi_{j}(t){\rm d}t=\sum_{r\in{\mathbb{Z}}}\int_{\mathbb{R}}f(x-t)e(\lambda\left[t\right]_{\mathsf{i}}^{c})\psi_{j(r,\lambda)}(t){\rm d}t\\ =\bigg(\sum_{r>0}+\sum_{r\leq 0}\bigg)\int_{\mathbb{R}}f(x-t)e(\lambda\left[t\right]_{\mathsf{i}}^{c})\psi_{j(r,\lambda)}(t){\rm d}t\coloneqq\mathsf{I}+\mathsf{II}\text{,}$
where $j(r,\lambda)$ is the unique $j\in\mathbb{Z}$ such that $2^{r}\leq|\lambda|^{\frac{1}{c}}2^{j}<2^{r+1}$, i.e.:
$j(r,\lambda)=\big\lfloor\log_{2}\big(2^{r}|\lambda|^{-\frac{1}{c}}\big)\big\rfloor=\big\lfloor r+\log_{2}\big(|\lambda|^{-\frac{1}{c}}\big)\big\rfloor=r+\big\lfloor\log_{2}\big(|\lambda|^{-\frac{1}{c}}\big)\big\rfloor\text{.}$
For the first summand in the last line of (6.13) we have the following pointwise bound
$|\mathsf{I}|\leq\sum_{r\in\mathbb{N}}\sup_{j\in\mathbb{Z}}\sup_{2^{r}\leq|\lambda|^{\frac{1}{c}}2^{j}\leq 2^{r+1}}\big|\mathcal{F}_{\mathbb{R}}^{-1}\big[\mathcal{F}_{\mathbb{R}}(f)(\cdot)H_{j}^{\mathsf{i}}(\cdot,\lambda)\big](x)\big|\text{,}$
and by Lemma 6.2, we get that
(6.14) $\|\mathsf{I}\|_{L^{2}(\mathbb{R})}\lesssim\sum_{r\in\mathbb{N}}(c-1)^{-\frac{1}{2}}2^{-\frac{r}{6}}\|f\|_{L^{2}(\mathbb{R})}\lesssim(c-1)^{-\frac{1}{2}}\|f\|_{L^{2}(\mathbb{R})}\text{.}$
It remains to treat the second summand, which we bound pointwise using the Hardy–Littlewood maximal function and maximal truncations of the Hilbert tranform. Firstly, note that
$\left|\mathsf{II}\right|=\sum_{r\leq 0}\int_{\mathbb{R}}f(x-t)\left(e(\lambda\left[t\right]_{\mathsf{i}}^{c})-1\right)\psi_{j(r,\lambda)}(t){\rm d}t+\sum_{r\leq 0}\int_{\mathbb{R}}f(x-t)\psi_{j(r,\lambda)}(t){\rm d}t\text{.}$
On the one hand, we have
$\Big|\sum_{r\leq 0}\int_{\mathbb{R}}f(x-t)\left(e(\lambda\left[t\right]_{\mathsf{i}}^{c})-1\right)\psi_{j(r,\lambda)}(t){\rm d}t\Big|\leq\sum_{r\leq 0}2^{rc}\mathrm{M}f(x)\lesssim\mathrm{M}f(x)\text{,}$
and on the other hand, we have
$\sum_{r\leq 0}\int_{\mathbb{R}}f(x-t)\psi_{j(r,\lambda)}(t){\rm d}t=\sum_{j\leq\lfloor\log_{2}(|\lambda|^{-\frac{1}{c}})\rfloor}\int_{\mathbb{R}}f(x-t)\psi_{j}(t){\rm d}t\leq\mathcal{H}^{*}f(x)\text{,}$
where $\mathcal{H}^{*}f(x)=\sup_{k\in{\mathbb{Z}}}\Big|\sum_{j\leq k}\int_{\mathbb{R}}f(x-t)\psi_{j}(t){\rm d}t\Big|$. Since both $\mathrm{M}$ and $\mathcal{H}^{*}$ are bounded operators on $L^{2}$, we immediately get that $\|\mathsf{II}\|_{L^{2}(\mathbb{R})}\lesssim\|f\|_{L^{2}(\mathbb{R})}$.
Taking the above estimate into account together with (6.13) and (6.14) concludes the proof. ∎